# HAT: Hybrid Attention Transformer for Image Restoration

### v2

@@ V-D Study on the pre-training strategy
From Tab. XI, we can see that HAT can benefit greatly from the pre-training strategy, by comparing the performance of HAT and HAT. To show the superiority of the proposed same-task pre-training, we also apply the multi-related-task pre-training [21] to HAT for comparison using full ImageNet, under the same training settings as  [21]. As depicted as Tab. X, the same-task pre-training performs better, not only in the pre-training stage but also in the fine-tuning process. From this perspective, multi-task pre-training probably impairs the restoration performance of the network on a specific degradation, while the same-task pre-training can maximize the performance gain brought by large-scale data. To further investigate the influences of our pre-training strategy for different networks, we apply our pre-training to four networks: SRResNet (1.5M), RRDBNet (16.7M), SwinIR (11.9M) and HAT (20.8M), as shown in Fig. 9. First, we can see that all four networks can benefit from pre-training, showing the effectiveness of the proposed same-task pre-training strategy. Second, for the same type of network (i.e., CNN or Transformer), the larger the network capacity, the more performance gain from pre-training. Third, although with less parameters, SwinIR obtains greater performance improvement from the pre-training compared to RRDBNet. It suggests that Transformer needs more data to exploit the potential of the model. HAT obtains the largest gain from pre-training, indicating the necessity of the pre-training strategy for such large models. Equipped with big models and large-scale data, we show that the performance upper bound of this task can be significantly extended.
VI Experimental Results
VI-A Training Settings
TABLE X: Quantitative results on PSNR(dB) of HAT using two kinds of pre-training strategies on $\times$4 SR under the same training setting.
Strategy Stage Set5 Set14 Urban100
Multi-related-task pre-training 32.94 29.17 28.05
pre-training fine-tuning 33.06 29.33 28.21
Same-task pre-training 33.02 29.20 28.11
pre-training(ours) fine-tuning 33.07 29.34 28.28
TABLE XI: Quantitative comparison with state-of-the-art methods for classic image super-resolution on benchmark datasets. The best and second best results are marked in bold and underline. $\dagger$” indicates that methods adopt pre-training strategy on ImageNet.
Method Scale Training Dataset Set5 [87] Set14 [88] BSD100 [89] Urban100 [90] Manga109 [91]
@@ VI-B Classic Image Super-Resolution
Figure 11: Comparison of LAM results between SwinIR and HAT.
Quantitative results. Tab. XI shows the quantitative comparison of our approach with the state-of-the-art methods: EDSR [28], RCAN [10], SAN [30], IGNN [36], HAN [31], NLSN [32], RCAN-it [86], as well as approaches using ImageNet pre-training, i.e., IPT [18] and EDT [21]. We can see that our method significantly outperforms the other methods on all five benchmark datasets. Concretely, with the same depth and width, HAT surpasses SwinIR by 0.48dBsimilar-$\sim$0.64dB on Urban100 and 0.34dBsimilar-$\sim$0.45dB on Manga109. When compared with the approaches using pre-training, HAT also has large performance gains of more than 0.5dB against EDT on Urban100 for all three scales. Besides, HAT equipped with pre-training outperforms SwinIR by a huge margin of up to 1dB on Urban100 for $\times$2 SR. Moreover, the large model HAT-L can even bring further improvement and greatly expands the performance upper bound of this task. HAT-S with fewer parameters and similar computation can also significantly outperforms the state-of-the-art method SwinIR. (Detailed computational complexity comparison can be found in Sec. V-C.) Note that the performance gaps are much larger on Urban100, as it contains more structured and self-repeated patterns that can provide more useful pixels for reconstruction when the utilized range of information is enlarged. All these results show the effectiveness of our method.
VI-C Real-world Image Super-Resolution
Quantitative results. Table XII presents a quantitative comparison of our method with state-of-the-art approaches: ESRGAN [12], BSRGAN [92], Real-ESRGAN [38], DASR [94], and SwinIR [22]. All models are GAN-based, and our HAT models are trained using the BSRGAN degradation model [92] (i.e., HAT-1) and Real-ESRGAN degradation model [38] (i.e., HAT-2). For compared methods, we utilize the officially released models and evaluate them on four datasets. Real-SR-cano [95] and Real-Nikon [95] feature real-world data pairs from specific cameras. AIM2019-val, from the AIM2019 challenge [96], is a synthetic dataset with realistic, unknown degradations [^2] We use paired data of the validation set from this challenge. The competition officials do not release the degradation model.. Additionally, we construct a synthetic dataset using 100 DIV2K valid [84] images with a high-order degradation model [38], following DASR [94]. We use PSNR and LPIPS [97] as metrics, with PSNR measuring fidelity and LPIPS assessing perceptual quality. Since all of the methods are GAN-based, the PSNR performance of different models on various datasets is not consistent. Nevertheless, similar PSNR values suggest comparable fidelity among the methods. Notably, HAT-1 achieves the best balance between PSNR and LPIPS among the three methods. HAT-2 obtains the best performance on all four datasets, indicating that it generates the results with the best perceptual quality.
TABLE XII: Quantitative comparison with state-of-the-art methods for real-world image super-resolution on benchmark datasets. The best and second best results are marked in bold and underline.
@@ VI-C Real-world Image Super-Resolution
Figure 12: Real-world image super-resolution results on SR ×$\times 4$. HAT-1 uses BSRGAN degradation model and HAT-2 adopts Real-ESRGAN degradation model.
VI-D Image Denoising
TABLE XIII: Quantitative comparison (average PSNR) with state-of-the-art methods for grayscale image denoising on benchmark datasets. The best nad second best results are marked in bold and underline, respectively.
 Dataset  $\sigma$ DnCNN [7] IRCNN [9] FFDNet [68] NLRN [39] RNAN [11] MWCNN [98] DRUNet [13] SwinIR [22] Restormer [20] SCUNet [74] HAT (ours)
Set12 [7] 15 32.86 32.76 32.75 33.16 - 33.15 33.25 33.36 33.42 33.43 33.49
25 30.44 30.37 30.43 30.80 - 30.79 30.94 31.01 31.08 31.09 31.13
50 27.18 27.12 27.32 27.64 27.70 27.74 27.90 27.91 28.00 28.04 28.07
BSD68 [89] 15 31.73 31.63 31.63 31.88 - 31.86 31.91 31.97 31.96 31.99 31.99
25 29.23 29.15 29.19 29.41 - 29.41 29.48 29.50 29.52 29.55 29.52
50 26.23 26.19 26.29 26.47 26.48 26.53 26.59 26.58 26.62 26.67 26.60
Urban100 [90] 15 32.64 32.46 32.40 33.45 - 33.17 33.44 33.70 33.79 33.88 33.99
25 29.95 29.80 29.90 30.94 - 30.66 31.11 31.30 31.46 31.58 31.67
50 26.26 26.22 26.50 27.49 27.65 27.42 27.96 27.98 28.29 28.56 28.62
TABLE XIV: Quantitative comparison (average PSNR) with state-of-the-art methods for color image denoising on benchmark datasets. The best, second best results are marked in bold and underline, respectively.
     Dataset  $\sigma$ DnCNN [7] IRCNN [9] FFDNet [68] RNAN [11] RDN [29] IPT [18] DRUNet [13] SwinIR [22] Restormer [20] SCUNet [74] HAT (ours)
CBSD68 [89] 15 33.90 33.86 33.87 - - - 34.30 34.42 34.40 34.40 34.42
@@ VI-D Image Denoising
TABLE XV: Quantitative comparison (average PSNR/SSIM) with state-of-the-art methods for JPEG compression artifacts reduction on benchmark datasets. The best and second best results are marked in bold and underline, respectively.
@@
VI-D1 Grayscale image denoising
Tab. XIII shows the quantitative comparison on grayscale image denoising of our approach with the state-of-the-art method: DnCNN [7], IRCNN [9], FFDNet [68], NLRN [39], MWCNN [98], DRUNet [13], SwinIR [22], Restormer [20] and SCUNet [74]. Compared to SwinIR [22], HAT achieves better performance on all datasets with multiple noise levels. On Urban100, HAT achieves the largest performance gain by up to 0.64dB for $\sigma$ 50. Compared to current state-of-the-art methods Restormer [20] and SCUNet [74], HAT can still outperform the former and obtains comparable performance with the latter.
@@ VI-D2 Color image denoising
Tab. XIV shows the quantitative comparison results on color image denoising of our approach with the state-of-the-art methods: DnCNN [7], IRCNN [9], FFDNet [68], RNAN [11], RDN [69], IPT [18], DRUNet [13], SwinIR [22], Restormer [20] and SCUNet [74]. We can observe that HAT achieves the best performance on almost all four benchmark datasets. Specifically, HAT outperforms SwinIR from 0.24dB to 0.38dB and surpasses SCUNet by a large margin of 0.19dB on Urban100 with $\sigma=15$.
We present the visual results of different methods in Fig. 14. For images “img 039”, “img 042” and “img 074” on Urban100, HAT reconstructs complete and clear edges, whereas other methods cannot produce complete lines or suffer from severe blurring effects. For the image “img 085” on Urban100, our method successfully restores the correct shape and clear texture, while other methods all fail. All these results demonstrate the superiority of the proposed HAT on image denoising.
@@ VI-E JPEG compression artifacts Reduction
Tab. XV shows the quantitative comparison results on JPEG compression artifacts reduction of our approach with the state-of-the-art methods: ARCNN [6], DnCNN [7], RNAN [11], DRUNet [13], FBCNN [101], SwinIR [22]. On Classic5 and LIVE1, HAT only achieves comparable performance to SwinIR. We consider this is because the demand for the model fitting ability of this task has approached saturation, particularly for low-resolution images. We further provide the performance comparison on Urban100. Then we can see that HAT achieves considerable performance gains over SwinIR, up to 0.18dB for JPEG quality $q=40$. This can be attributed to the presence of a large number of regular textures and repeating patterns in the images of the Urban100 dataset. HAT is capable of activating more pixels for restoration. With a larger receptive field, it can restore sharper edges and textures.

### v3

@@ V-D Study on the Pre-training Strategy
From Tab. IX, we can see that HAT can benefit greatly from the pre-training strategy, by comparing the performance of HAT and HAT. To show the superiority of the proposed same-task pre-training, we also apply the multi-related-task pre-training [21] to HAT for comparison using full ImageNet, under the same training settings as  [21]. As depicted as Tab. VIII, the same-task pre-training performs better, not only in the pre-training stage but also in the fine-tuning process. From this perspective, multi-task pre-training probably impairs the restoration performance of the network on a specific degradation, while the same-task pre-training can maximize the performance gain brought by large-scale data. To further investigate the influences of our pre-training strategy for different networks, we apply our pre-training to four networks: SRResNet (1.5M), RRDBNet (16.7M), SwinIR (11.9M) and HAT (20.8M), as shown in Fig. 9. First, we can see that all four networks can benefit from pre-training, showing the effectiveness of the proposed same-task pre-training strategy. Second, for the same type of network (i.e., CNN or Transformer), the larger the network capacity, the more performance gain from pre-training. Third, although with less parameters, SwinIR obtains greater performance improvement from the pre-training compared to RRDBNet. It suggests that Transformer needs more data to exploit the potential of the model. HAT obtains the largest gain from pre-training, indicating the necessity of the pre-training strategy for such large models. Equipped with big models and large-scale data, we show that the performance upper bound of this task can be significantly extended.
TABLE IX: Quantitative comparison with state-of-the-art methods for classic image super-resolution on benchmark datasets. The best and second best results are marked in bold and underline. $\dagger$” indicates that methods adopt pre-training strategy on ImageNet.
Method Scale
Training
Dataset
Set5 [88] Set14 [89] BSD100 [90] Urban100 [91] Manga109 [92]
@@
Note: For fair comparison, GRL’s SSIM values are recalculated using BasicSR based on its official results, as the original reported values differ significantly.
VI Experimental Results
VI-A Training Settings
@@ VI-B Classic Image Super-Resolution
[…]
TABLE X: Quantitative comparison with state-of-the-art methods for real-world image super-resolution on benchmark datasets. The best and second best results are marked in bold and underline.
@@
Figure 12: Real-world image super-resolution results on SR $\times 4$. HAT-1 uses BSRGAN degradation model and HAT-2 adopts Real-ESRGAN degradation model.
VI-C Real-world Image Super-Resolution
Quantitative results. Table X presents a quantitative comparison of our method with state-of-the-art approaches: ESRGAN [12], BSRGAN [94], Real-ESRGAN [38], DASR [98], and SwinIR [22]. All models are GAN-based, and our HAT models are trained using the BSRGAN degradation model [94] (i.e., HAT-1) and Real-ESRGAN degradation model [38] (i.e., HAT-2). For compared methods, we utilize the officially released models and evaluate them on four datasets. Real-SR-cano [96] and Real-Nikon [96] feature real-world data pairs from specific cameras. AIM2019-val, from the AIM2019 challenge [97], is a synthetic dataset with realistic, unknown degradations [^2] We use paired data of the validation set from this challenge. The competition officials do not release the degradation model.. Additionally, we construct a synthetic dataset using 100 DIV2K valid [85] images with a high-order degradation model [38], following DASR [98]. We use PSNR and LPIPS [99] as metrics, with PSNR measuring fidelity and LPIPS assessing perceptual quality. Since all of the methods are GAN-based, the PSNR performance of different models on various datasets is not consistent. Nevertheless, similar PSNR values suggest comparable fidelity among the methods. Notably, HAT-1 achieves the best balance between PSNR and LPIPS among the three methods. HAT-2 obtains the best performance on all four datasets, indicating that it generates the results with the best perceptual quality.
TABLE XI: Quantitative comparison (average PSNR) with state-of-the-art methods for grayscale image denoising on benchmark datasets. The best nad second best results are marked in bold and underline, respectively.
 Dataset  $\sigma$ DnCNN [7] IRCNN [9] FFDNet [69] NLRN [40] MWCNN [100] DRUNet [13] SwinIR [22] Restormer [20] SCUNet [75] GRL [93] HAT (ours)
Set12 [7] 15 32.86 32.76 32.75 33.16 33.15 33.25 33.36 33.42 33.43 33.47 33.49
25 30.44 30.37 30.43 30.80 30.79 30.94 31.01 31.08 31.09 31.12 31.13
50 27.18 27.12 27.32 27.64 27.74 27.90 27.91 28.00 28.04 28.03 28.07
BSD68 [90] 15 31.73 31.63 31.63 31.88 31.86 31.91 31.97 31.96 31.99 32.00 31.99
25 29.23 29.15 29.19 29.41 29.41 29.48 29.50 29.52 29.55 29.54 29.52
50 26.23 26.19 26.29 26.47 26.53 26.59 26.58 26.62 26.67 26.60 26.60
Urban100 [91] 15 32.64 32.46 32.40 33.45 33.17 33.44 33.70 33.79 33.88 34.09 33.99
25 29.95 29.80 29.90 30.94 30.66 31.11 31.30 31.46 31.58 31.80 31.67
50 26.26 26.22 26.50 27.49 27.42 27.96 27.98 28.29 28.56 28.59 28.62
TABLE XII: Quantitative comparison (average PSNR) with state-of-the-art methods for color image denoising on benchmark datasets. The best, second best results are marked in bold and underline, respectively.
     Dataset  $\sigma$
DnCNN
[7]
IRCNN
[9]
FFDNet
[69]
RNAN
[11]
RDN
[29]
IPT
[18]
DRUNet
[13]
SwinIR
[22]
Restormer
[20]
SCUNet
[75]
HAT
CBSD68
[90]
15 33.90 33.86 33.87 - - - 34.30 34.42 34.40 34.40 34.42
@@ VI-C Real-world Image Super-Resolution
TABLE XIII: Quantitative comparison (average PSNR/SSIM) with state-of-the-art methods for JPEG compression artifacts reduction on benchmark datasets. The best and second best results are marked in bold and underline, respectively.
@@
Figure 16: Real-world image denoising results on the SIDD dataset [106].
VI-D Image Denoising
VI-D1 Grayscale Image Denoising
Tab. XI shows the quantitative comparison on grayscale image denoising of our approach with the state-of-the-art method: DnCNN [7], IRCNN [9], FFDNet [69], NLRN [40], MWCNN [100], DRUNet [13], SwinIR [22], Restormer [20] , GRL [93] and SCUNet [75]. Compared to SwinIR [22], HAT achieves better performance on all datasets with multiple noise levels. On Urban100, HAT achieves the largest performance gain by up to 0.64dB for $\sigma$ 50. Compared to current state-of-the-art methods Restormer [20] and SCUNet [75], HAT can still outperform the former and obtains comparable performance with the latter.
@@ VI-D2 Color Image Denoising
Tab. XII shows the quantitative comparison results on color image denoising of our approach with the state-of-the-art methods: DnCNN [7], IRCNN [9], FFDNet [69], RNAN [11], RDN [70], IPT [18], DRUNet [13], SwinIR [22], Restormer [20] and SCUNet [75]. We can observe that HAT achieves the best performance on almost all four benchmark datasets. Specifically, HAT outperforms SwinIR from 0.24dB to 0.38dB and surpasses SCUNet by a large margin of 0.19dB on Urban100 with $\sigma=15$.
We present the visual results of different methods in Fig. 14. For images “img 039”, “img 042” and “img 074” on Urban100, HAT reconstructs complete and clear edges, whereas other methods cannot produce complete lines or suffer from severe blurring effects. For the image “img 085”, our method successfully restores the correct shape and clear texture, while other methods all fail. All these results demonstrate the superiority of the proposed HAT on image denoising.
VI-D3 Real-world Image Denoising
We also evaluate our proposed method for real-world image denoising based on the SIDD dataset [106]. The compared methods include MPRNet [72], UFormer [19], MAXIM [67], HINet [107], Restormer [20], and NAFNet [73]. The quantitative and qualitative results are shown in Tab. 16 and Fig. XIV. Our method achieves the highest PSNR/SSIM scores and clearest results, demonstrating superior performance in real-world scenarios.
TABLE XIV: Real-world image denoising results on the SIDD dataset [106].
Method MPRNet [72] UFormer [19] MAXIM [67] HINet [107] Restormer [20] NAFNet [73] HAT (ours)
PSNR(dB)/SSIM 39.71/0.958 39.89/0.960 39.96/0.960 39.99/0.958 40.02/0.960 40.30/0.962 40.50/0.964
TABLE XV: Image deblurring results on the Gopro dataset [71].
Method MPRNet [72] HINet [107] MAXIM [67] Restormer [20] UFormer [19] NAFNet [73] GRL [93] HAT (ours)
PSNR(dB)/SSIM 32.66/0.959 32.71/0.959 32.86/0.961 32.92/0.961 32.97/0.967 33.69/0.967 33.93/0.968 33.96/0.968
Figure 17: Image deblurring results on the Gopro dataset [71].
@@ VI-E JPEG Compression Artifacts Reduction
Tab. XIII shows the quantitative comparison results on JPEG compression artifacts reduction of our approach with the state-of-the-art methods: ARCNN [6], DnCNN [7], RNAN [11], DRUNet [13], FBCNN [103], SwinIR [22]. On Classic5 and LIVE1, HAT only achieves comparable performance to SwinIR. We consider this is because the demand for the model fitting ability of this task has approached saturation, particularly for low-resolution images. We further provide the performance comparison on Urban100. Then we can see that HAT achieves considerable performance gains over SwinIR, up to 0.18dB for JPEG quality $q=40$. This can be attributed to the presence of a large number of regular textures and repeating patterns in the images of the Urban100 dataset. HAT is capable of activating more pixels for restoration. With a larger receptive field, it can restore sharper edges and textures.
@@
VI-F Image Deblurring
We conduct experiments to further validate the effectiveness of our method on image deblurring. Specifically, we evaluate several representative deblurring methods on the GoPro dataset [71]. The compared methods include MPRNet [72], HINet [72], MAXIM [67], Restormer [20], Uformer [19], NAFNet [73], and GRL [93]. As shown in Tab. XV, our HAT achieves the best performance among all compared approaches, with a PSNR of 33.96 dB. In addition, as illustrated in Fig. 17, HAT successfully reconstructs sharp and clear human silhouettes and portraits, recovering visually plausible details from heavily blurred inputs. These results confirm that our method not only achieves state-of-the-art performance in tasks such as SR, denoising, and compression artifact reduction, but also generalizes effectively to image deblurring. This further validates the superiority of our design.
@@ VII Conclusion
Acknowledgments
This work was supported in part by National Natural Science Foundation of China (Grant No.62276251, 62272450); in part by Macau Science and Technology Development Fund under 001/2024/SKL, 0022/2022/A1, and 0119/2024/RIB2; in part by Research Committee at University of Macau under MYRG-GRG2023-00058-FST-UMDF; in part by the Guangdong Basic and Applied Basic Research Foundation under Grant 2024A1515012536.